OpenStudy (anonymous):

is (x+2) a factor of 3x^2-x-14?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If x + 2 is the factor of 3x^2-x-14, then x = -2 is its zero. Try by replace x by -2, u've got 3(-2)^2-(-2)-14 = 3*4 + 2 - 14 = 0 (Right) => Yes, it is

OpenStudy (anonymous):

The two factors of the equations are (x+2) and (3x-7), so YES